HUBBLE SPACE TELESCOPE DAILY REPORT #1100 PERIOD COVERED: 0700 EST Daily Status Report as of 081/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 4709 (WF/PC-2 Internal Monitor) The WF/PC camera was used to make a series of routine ultraviolet flat calibration images to monitor the instrument's performance. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were noted. 1.2 Completed FOS/BL 5619 (SMOV FGS/FOS Fine Alignment (Small Apertures) Revitalized) The Faint Object Spectrograph (Blue Side) and the Fine Guidance Sensors were used to make precise V2-V3 position and size alignment calibration observations using standard star STAR-22515. A large series of observations were received and displayed. For two of the observations the dwell scan pattern was not near the target. No other anomalies were apparent. Much of the data still remains on the science and engineering tape recorders, and will be analyzed later. 1.3 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and for contamination. The observations were monitored by the OSS, and the data quality was assessed. No anomalies were noted. 1.4 Completed WF/PC-2 5612 (SMOV WF/PC-2 Decontamination) The WF/PC-2 was used to perform decontamination operations. Observations of standard A+81D266, and K-Spot observations were also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled :7 Successful :7 2.3 Significant HSTARs: None 2.4 Operations Notes: No operational problems were encountered during the past 24 hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WF/PC-2 4709 (WF/PC-2 Internal Monitor) 3.2 Complete WF/PC-2 5109 (A Deep Survey at High Galactic Latitudes(WC06): Cycle 4) 3.3 Complete Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3. Darks) 3.4 Complete Two Sets of WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) 3.5 Complete Two sets of WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Large Set) 3.6 Complete GHRS/1 5544 (GHRS Side One Spatial PSF Characterization: Cycle 4) 3.7 Complete WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-